OpenStudy (anonymous):

Find sin θ if cot θ = - 2 and cos θ < 0. I need help I give medals the options are -1/2 -5 sqrt5/2 sqrt5/5

OpenStudy (anonymous):

@rob1525 ?

OpenStudy (anonymous):

Hey, I answered on the other question.

OpenStudy (rob1525):

well -2 is not in the form of radians. I'm not sure yet.

OpenStudy (anonymous):

@ivanmlerner so what is the answer to this one sorry I didnt put the options on that one ...ill give a medal though

OpenStudy (anonymous):

No problem, I took some time to answer its ok. But I dont see my answer on your options. If you look at the picture I got 1/sqrt(5)

OpenStudy (rob1525):

so this picture shows the cot theta = -2/1 now do the sin of the same angle. do you know what i mean?

OpenStudy (anonymous):

Oooh I see now, 1/sqrt(5)=sqrt(5)/5 and thats one of the options

OpenStudy (rob1525):

Actually the picture is wrong because 1 is positive so its actully in the 2 quadrant.

OpenStudy (anonymous):

Nice way @rob1525.

OpenStudy (anonymous):

@ivanmlerner can you help with : Factor the algebraic expression below in terms of a single trigonometric function. sin 2x + sin x - 2

OpenStudy (anonymous):

Sure, do you know the trig identity for sin(2x)? This question has something to do with that but it is just something extra, do you have a basic understanding of complex numbers?

OpenStudy (anonymous):

is it 2sin(x)cos(x) @ivanmlerner

OpenStudy (anonymous):

Yes, so you just need to use that, and the formula cos=sqrt(1-sin^2) to transform everything into sin

OpenStudy (anonymous):

no sure how to though thats as far as i know ... can you run through it for me

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

@ivanmlerner

OpenStudy (anonymous):

Sure, you have\[\sin(2x)+\sin x-2\]Using the formula for sin(2x)\[2\sin x \cos x+\sin x-2\]What the problem wants is that you use only one trig function (sin or cos in this case) in the formula, so we need to get rid of the cos. and for that we use the formula sin^2+cos^2=1 and this formula can be rearanged to cos=sqrt(1-sin^2) so we need only to substitute now on the formula:\[2\sin x \cos x+\sin x-2=2\sin x \sqrt{1-\sin^2 x}+\sin x -2\]And I believe thats it because we only use sin(x) in this formula.

OpenStudy (anonymous):

sqrt5/5